OpenStudy (anonymous):

Reduction formula ∫sin^7x dx

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

@ajprincess

ganeshie8 (ganeshie8):

put cosx = t -sinx dx = dt

ganeshie8 (ganeshie8):

\(\large \mathbb{\int \sin^7x dx }\) \(\large \mathbb{\int \sin^6x ~ \sin xdx }\) \(\large \mathbb{\int (\sin^2x)^3 ~ \sin xdx }\) \(\large \mathbb{\int (1-\cos^2x)^3 ~ \sin xdx }\) put cosx = t -sinx dx = dt \(\large \mathbb{-\int (1-t^2)^3 ~ dt}\)

ganeshie8 (ganeshie8):

expand the cubic and integrate each term

OpenStudy (anonymous):

the answer is -cosx+cos^3x-3cos^5x/5+cos^7x/7+c @ganeshie8

OpenStudy (anonymous):

right?

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\color{red}{\checkmark}\)

OpenStudy (anonymous):

Thank you sir :)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

:)